### v1

# ClaraVid: A Holistic Scene Reconstruction Benchmark From Aerial Perspective With Delentropy-Based Complexity Profiling

Abstract
The development of aerial holistic scene understanding algorithms is hindered by the scarcity of comprehensive datasets that enable both semantic and geometric reconstruction. While synthetic datasets offer an alternative, existing options exhibit task-specific limitations, unrealistic scene compositions, and rendering artifacts that compromise real-world applicability. We introduce ClaraVid, a synthetic aerial dataset specifically designed to overcome these limitations. Comprising 16,917 high-resolution images captured at 4032×3024 from multiple viewpoints across diverse landscapes, ClaraVid provides dense depth maps, panoptic segmentation, sparse point clouds, and dynamic object masks, while mitigating common rendering artifacts. To further advance neural reconstruction, we introduce the Delentropic Scene Profile (DSP), a novel complexity metric derived from differential entropy analysis, designed to quantitatively assess scene difficulty and inform reconstruction tasks. Utilizing DSP, we systematically benchmark neural reconstruction methods, uncovering a consistent, measurable correlation between scene complexity and reconstruction accuracy. Empirical results indicate that higher delentropy strongly correlates with increased reconstruction errors, validating DSP as a reliable complexity prior. Currently under review, upon acceptance the data and code will be available at rdbch.github.io/ClaraVid.
{strip}[Uncaptioned image]
Figure 1: Overview of ClaraVid: (a) Multi-viewpoint UAV acquisition: High-resolution aerial imagery is captured from multiple altitudes, ensuring diverse perspectives. (b) High-fidelity, diverse environments: complex urban, suburban, and natural landscapes. (c) Multimodal ground truth: Pixel-level and scene-level multimodal data for holistic scene reconstruction and semantic mapping.
1 Introduction
Semantic and geometric scene understanding derived from low-altitude uncrewed aerial vehicle (UAV) imagery has become increasingly essential for applications such as disaster response  [1], urban mapping [2], and infrastructure inspection [3], where accurate and high-fidelity environmental perception directly informs critical decision-making processes. Recent advancements in neural reconstruction techniques, notably Neural Radiance Fields (NeRFs)[4, 5, 6] and Gaussian Splatting approaches[7, 8, 9], have demonstrated significant potential in enabling scene-level querying and novel-view synthesis, offering a transformative approach to reconstructing complex 3D environments from sparse 2D observations. However, evaluating holistic scene mapping in urban environments—encompassing both semantic and geometric aspects—remains constrained by the limitations of existing datasets. Real-world benchmarks providing comprehensive multi-modal ground truth for the joint tasks are inexistent for urban scenarios, and synthetic alternatives suffer from critical shortcomings: they often lack key sensing modalities such as dense semantic segmentation [10, 11], are not collected for scene mapping applications [12, 13], or are captured from restricted viewpoints that hinder a full assessment of reconstruction quality and generalization [14].
In this work, we introduce ClaraVid, a synthetic aerial dataset crafted to support the training and evaluation of holistic scene representations while also enabling lower-level perception tasks. ClaraVid employs a mapping-centric approach, emulating UAV mapping missions to reflect real-world aerial survey workflows. Spanning 5 distinct environments across 8 missions, the dataset provides high-resolution imagery at 4032 × 3024 pixels, collected simultaneously from 3 different viewpoints, accompanied by multimodal annotations including dense depth maps, panoptic segmentation for buildings and actors, and dynamic object masks to precisely delineate moving entities—a distinctive enhancement over prior synthetic datasets. Furthermore, it includes multi-resolution multi-modal scene-level point clouds approximating LiDAR scans, broadening its utility for geometric analysis. Data quality is ensured through an optimized rendering pipeline that addresses challenges such as foliage culling and level-of-detail inconsistencies, resulting in enhanced fidelity for scene reconstruction and semantic segmentation. The dataset is complemented by a benchmarking framework designed to assess semantic and geometric reconstruction accuracy, while also supporting frame-level tasks such as semantic segmentation. Relevant image samples are presented in ClaraVid: A Holistic Scene Reconstruction Benchmark From Aerial Perspective With Delentropy-Based Complexity Profiling.
A major challenge in UAV-based scene reconstruction is quantifying scene complexity in a way that correlates with the difficulty of downstream reconstruction tasks. Currently the base solution is based on an heuristic assessments rather than formalized metrics, making it difficult to systematically predict the impact of scene structure on the reconstruction. Delentropy[15] , an entropy measure derived from gradient field analysis, has previously been employed to analyze information content in pixel-level tasks [16, 17, 18]. In this work, we adapt delentropy for scene profiling, and propose a novel Delentropic Scene Profile, applied in the context of neural reconstruction, leveraging its ability to capture the spatial organization of scene features through second-order entropy analysis. We empirically demonstrate a strong correlation between the proposed metric and the final reconstruction error across multiple approaches.
To summarize, our paper introduces the following primary contributions to advance UAV-based holistic scene understanding: (1) We present ClaraVid, a novel synthetic aerial dataset specifically designed for holistic scene reconstruction, comprising high-resolution imagery, multi-modal annotations (dense depth maps, panoptic segmentation, sparse point clouds, and dynamic object masks), and realistic rendering tailored to aerial mapping applications. (2) We introduce the Delentropic Scene Profile (DSP), a delentropy-based metric for quantitatively assessing scene complexity. (3) We empirically validate DSP’s utility, demonstrating strong correlations between scene complexity and reconstruction performance, offering a principled framework for guiding method selection and optimizing neural reconstruction pipelines.
Table 1: Comparison of related datasets from aerial and automotive domains relevant for holistic scene representation.
Dataset Year Type Camera Calib. Seg. Depth Scene PCL Dynamic Mask For Mapping Viewpoints Render Issues Image Size Samples
Automotive Datasets
A2D2 [19] 2020 R Panoptic - - - 6 - 1440x960 41.2k
Waymo [20] 2020 R Panoptic - - - 5 - 1920x1280 390k
Urban Syn [21] 2023 S Panoptic - - - 1 Low 2048x1024 7.5k
Aerial Datasets
UDD [22] 2018 R Semantic 1 - 4096x2160 0.3k
Aeroscapes [23] 2019 R Semantic 1 - 1280x720 3.2k
UAVid [24] 2020 R Semantic 1 - 4096x2160 0.4k
UrbanScene3D [11] 2022 S/R Instance* RGB/Ins 1 - 8192x5460 128k
Ommo [25] 2023 R RGB 1 - 4096x2160 14.7k
Matrix City [10] 2023 S RGB 1 Low* 1920x1080 410k
Syndrone [12] 2023 S Semantic 3 Medium 1920x1080 72k
DDos [14] 2024 S Semantic 1 Medium 1280x720 34k
ClaraVid (ours) - S Panoptic RGB/Sem/Ins 3 Low* 4032x3024 16.3k
2 Related work
Aerial datasets The progress in semantic and geometric scene understanding has been largely driven by multi-modal automotive datasets, real [26, 19, 20] and synthetic [27, 28, 21]. These datasets provide strong benchmarks that integrate LiDAR, stereo imagery, and detailed annotations for tasks enabling autonomous navigation. In contrast, real-world aerial datasets, are often small in scale and fragmented in scope: semantic segmetation[24, 29, 22, 30], depth estimation [31, 32], object tracking [33], etc. Synthetic aerial datasets, created using graphics engines, offer a scalable alternative with rich semantic content. However, they come with significant limitations. Their resolution is typically limited to 2K [13], whereas real datasets can reach 4K [24, 22, 29]. Rendering artifacts, such as issues with foliage rendering [34] and inconsistent levels of detail [13, 12], reduce their realism. Additionally, key modalities like semantic segmentation are frequently missing [21, 10], making them less effective for mapping applications. Claravid addresses these limitations directly by providing a dataset specifically designed for mapping tasks. It includes multi-view imagery, panoptic annotations, and dense depth maps, offering a high-fidelity benchmark targeted towards real-world applications. A summarized overview of the presented datasets is provided in Table 1.
Neural scene representation of urban scenarios from aerial perspective Recent advancements in neural scene representations have established robust frameworks for modeling complex, real-world environments. NeRFs, introduced by Mildenhall etal[4], have proven highly effective across a range of tasks–where further extensions–shown promising results from novel view synthesis [35, 36] to surface reconstruction [37, 38, 39, 40] and panoptic reconstruction [41, 42]. In parallel, Gaussian Splatting [7] has emerged as a particle-based alternative that bypasses the computational overhead of neural networks, thereby enabling significantly accelerated training and rendering for real-time, large-scale scene representations [7, 43, 8, 44]. However, many of these approaches are relying on prior data for semantic information that can be obtained typically from GT data or as pseudolabels. Finally, while both paradigms have been applied to urban scene reconstruction from aerial perspectives [45, 46], they largely omit semantic information—primarily due to the absence of comprehensive datasets. We introduce our dataset specifically designed to benchmark these diverse methodologies within a unified framework that integrates all relevant modalities from a low aerial viewpoint.
Dataset complexity through the information-theory lens Differential entropy has emerged as a robust information-theoretic tool for quantifying image complexity, building on Shannon’s definition of entropy[47] as a measure of uncertainty in data. Larkin’s formulation of delentropy [15] exemplifies this approach, extending traditional entropy to incorporate spatial structure by using a gradient-based joint density (deldensity) that captures pixel co-occurrence and image detail. This enables a more comprehensive complexity measure than conventional metrics: whereas basic Shannon entropy ignores spatial relationships among pixels and GLCM-based texture entropy[48] captures only limited local patterns, delntropy integrates both local and global features of an image’s intensity distribution. Rahane et al.[16] compared large-scale image datasets using entropy distributions (pixel-level, texture, and delentropy), finding that higher-order measures like delentropy align well with dataset learning difficulty and correlate with known variations in model performance for the semantic segmentation task. Similarly, Cagas et al. [18] observed that a dataset’s complexity distribution—quantified via delentropy—correlates with generative model fidelity, informing how increasing image complexity can demand larger training sets for GANs[49]. We leverage delentropy to construct a scene-level complexity measure that serves as a proxy for assessing reconstruction difficulty in aerial missions.
3 Scene complexity through delentropy
3.1 Theoretical background
Image complexity can be defined as the variety of features and details within an image. Traditional entropy is a foundational abstraction in information theory introduced by Shannon[47]. It is used as a measure of uncertainty in data, specifically the variation in the distribution of pixel intensities of an image in grayscale format. To formalize delentropy mathematically, we start with its foundation in the gradient field of an image. Given an image $I$, the gradient components are computed as the partial derivatives:
$f_{x}=\frac{\partial I}{\partial x},\quad f_{y}=\frac{\partial I}{\partial y}$ (1)
These components represent the rate of change in the $x$ (vertical) and $y$ (horizontal) directions, respectively, capturing edges and variations. In practice, these are computed using the Sobel operator[50]. The joint probability density function, the deldensity $p(f_{x},f_{y})$ is estimated using a 2D histogram of $f_{x}$ and $f_{y}$, formally denoted using the Kronecker delta $\delta$ as follows:
$p_{i,j}=\frac{1}{4XY}\sum_{n=-N}^{N-1}\sum_{m=-M}^{M-1}\delta_{i,f_{x}(m,n)}\delta_{j,f_{y}(m,n)}$ (2)
where $p_{ij}$ represents the joint probability of observing the gradient values (i,j) at any pixel, $N,M$ define the spatial window size over which the gradient distribution is computed and $X,Y$, the image width and height. Further the sum over all non-zero gradient value pairs is normalized such that it represents a probability density function $\sum_{i,j}p_{ij}=1$. Delentropy[15], denoted as $H_{\text{del}}$, is then defined as the joint entropy of this distribution. If $p_{ij}$ is the probability of $f_{x}$ and $f_{y}$ falling into the $i$-th and $j$-th bins, respectively, the discrete form is:
$H_{\text{del}}=-\frac{1}{2}\sum_{i}^{I}\sum_{j}^{J}p_{ij}\log_{2}p_{ij}$ (3)
such that $I$ and $J$ represent the number of bins (discrete cells) in the 2D distribution, while the factor $\tfrac{1}{2}$ is derived from  [51] and is maintained in our case to preserve the scale consistency without significantly affecting the metrics. Further we will denote $H_{del}(I)$ the delentropy of image $I$.
Figure 2: Delentropic Scene Profile of Mill19 Building[46]. The histogram represents the delentropy distribution, with overlaid sample images illustrating complexity variations. The dashed curve denotes the fitted Beta distributionand the vertical red line marks the mean delentropy.
3.2 Delentropic Scene Profile
The delentropy measure, $H_{\text{del}}$, quantifies structural complexity for an individual image. To extend this to a collection of images at scene level, we introduce the novel concept of Delentropic Scene Profile ($DSP_{S}$), a quantitative framework capturing the complexity of a multi-view scene and its implications for reconstruction accuracy. A scene $S$ is a finite set of images, $S=\{I_{1},I_{2},\ldots,I_{N}\}$, where each $I_{k},k=1,\ldots,N$ corresponds to a unique viewpoint of a shared environment, collectively encoding its 3D structure. The delentropy set of $S$ is defined as $\{H_{\text{del},k}\}_{k=1}^{N}$, with $H_{\text{del},k}=H_{\text{del}}(I_{k})$ computed over the grayscale pixels relevant to reconstruction, assigning zero informational content to excluded regions (e.g., masked dynamic objects) to preserve measure integrity.
The Delentropic Scene Profile, $\text{DSP}_{S}$, is the statistical distribution of $\{H_{\text{del},k}\}$, explicitly aproximated to a truncated Beta distribution parameterized by $(\alpha,\beta,a,b)$:
$\text{DSP}_{S}$ $=Beta(H_{\text{del}}\mid\alpha,\beta,a,b)$
$=\frac{(H_{\text{del}}-a)^{\alpha-1}(b-H_{\text{del}})^{\beta-1}}{(b-a)^{\alpha+\beta-1}B(\alpha,\beta)},$ (4)
[…]
$L(\alpha,\beta,a,b)=\prod_{k=1}^{N}f(H_{\text{del},k}\mid\alpha,\beta,a,b).$ (5)
In the absence of closed-form solutions, we first estimate $\alpha$ and $\beta$ using the Method of Moments (MoM), with the sample moments appropriately adjusted for the support $[a,b]$. These MoM estimates then serve as initial values for the Newton-Raphson algorithm, which iteratively refines the parameters to maximize the likelihood function.
Figure 3: ClaraVid Delentropic Scene Profiles as the distribution of delentropy scores across scenes. Each density plot represents delentropy values per image, with colors indicating the most frequent semantic class. The red horizontal line marks the mean delentropy, while the dashed curve shows the fitted Beta distribution. Water-dominated scenes exhibit long lower tails due to low-texture regions, whereas higher delentropy values, reflecting greater visual complexity, are prevalent in dense vegetation (trees, wheat fields, etc). Statistical parameters ($\mu$, $\sigma$, $\alpha$, $\beta$) quantify scene complexity.
3.3 Motivation for Beta Distribution
The choice of distributional model for $\text{DSP}_{S}$ was guided by the analysis of delentropy profiles of diverse datasets—ClaraVid, Skydrone[12], Skyscapes[30], UAVid[24], 3D Matrix City [10], Mill19 [46], and UrbanScene3D [11]—observing varied distributional forms contingent on the reconstruction context. A Gaussian distribution, while apt for uniformly complex scenes, fails to capture skewed or long-tailed distributions prevalent in heterogeneous environments (for example where we have a body of water or sky regions). Alternatives such as the Pareto or Gamma distributions accommodate skewness but lack flexibility for U-shaped profiles, which emerge in scenarios with pronounced bimodal complexity (e.g., a field with ocassional high-detail trees). The Beta distribution, however, adeptly models this spectrum—uniform, skewed, and U-shaped—owing to its shape parameters $\alpha$ and $\beta$, requiring only four parameters (including bounds $a$ and $b$) for a parsimonious yet expressive fit. This simplicity affords intuitive interpretation: $\alpha$ and $\beta$ directly govern the density’s form, while $a$ and $b$ anchor it to the empirical delentropy range. Multimodal distributions were eschewed, as their complexity—necessitating mixture models with additional parameters—hampers interpretability, risks overfitting, particularly given the finite sample sizes of $\{H_{\text{del},k}\}$ and did not show any improvements as compared with the best one. The Beta distribution emerges as the best choice, balancing flexibility, robustness, and theoretical clarity.
3.4 Delentropic Scene Profile Interpretation
The Delentropic Scene Profile (DSP) quantifies scene complexity through three key statistical descriptors: the mean ($\mu$), standard deviation ($\sigma$), and the shape parameters ($\alpha,\beta$) of the Beta-distributed delentropy values, $H_{\text{del}}\sim\text{Beta}(\alpha,\beta,a,b)$. The mean $\mu$ encapsulates the overall structural complexity of a scene. Low-complexity environments ($\mu<2.5$) predominantly contain large uniform regions, such as clear skies, bodies of water, or uniformly textured surfaces, whereas high-complexity scenes ($\mu>4.75$) are characterized by high frequency textures such as dense vegetation, roof tilings, etc. The standard deviation $\sigma$ quantifies the dispersion of delentropy values across the scene, with higher values indicating greater structural heterogeneity. The shape parameters $\alpha$ and $\beta$ define the skewness of the delentropy distribution: $\alpha<\beta$ signifies a predominance of low-detail imagery, while $\alpha>\beta$ indicates a tendency towards high-complexity views. When $\alpha\approx\beta<1$, the distribution is bimodal, capturing scenes with both extreme simplicity and complexity, whereas $\alpha\approx\beta>1$ suggests a concentration of delentropy values around moderate complexity levels. Together, these descriptors provide a quantitative framework for characterizing aerial scene complexity, facilitating the selection of optimal reconstruction strategies. A visual example is presented in Figure 2.
Discussion In the context of neural scene reconstruction, the DSP provides a lens to interpret different types of predictive uncertainty—epistemic (coverage gaps in training or inherent reconstruction method constraints) and heteroscedastic (unreducible errors). A long tail of low-delentropy images (e.g., large uniform regions such as sky or water) offers minimal geometric cues, intensifying epistemic and model uncertainty. Here, standard strategies like increasing the number of training samples or model size often fail to reduce error; instead, domain-specific solutions—such as incorporating depth priors have the potential to be more effective[52]. Conversely, a long tail of high-delentropy images, images with intricate textures and geometry, (i.e densely tiled roofs, trees, etc) can drive large variations when the model is under-parameterized, and additional viewpoints or higher network capacity to capture subtle details. In both extremes cases of delentropy values, heteroscedastic uncertainty limits how much the error can be reduced solely through oversampling or upscaled architectures. By revealing how complexity is distributed across scenes, the DSP informs targeted resource allocation and methodology choices, that ultimately can improve reconstruction fidelity.
4 ClaraVid Dataset
In this section we describe ClaraVid, a synthetic dataset designed for semantic and geometric reconstruction tasks. It comprises 16,917 multimodal frames collected across 5 diverse environments—urban, urban high, rural, highway, and natural landscapes— through 8 distinct UAV missions. The dataset is structured to serve as a benchmark for aerial perception tasks, such holistic scene reconstruction and the downstream supporting tasks such as panoptic segmentation, depth estimation, etc.
To ensure comprehensive scene coverage, ClaraVid employs a structured grid-based acquisition strategy, capturing images with an 80–90% overlap to balance redundancy and spatial continuity. Each mission is recorded from 3 distinct viewpoints, simulating a multi-UAV coordination paradigm. Data is acquired at varying altitudes and pitch angles: low (45–55m) 45°, mid (55–65m) 55°, and high (65–75m) 90°, with altitude adjustments dynamically adapting to terrain complexity. Importantly, the dataset does not consist of continuous image sequences; instead, there is an intentional 2–3 second pause between consecutive captures to simulate real-world aerial mapping scenarios, where UAVs acquire individual high-quality frames rather than video streams. ClaraVid spans a total area of 1.8 km2, with an average scene coverage of 0.22 km2, ensuring a rich variety of aerial perspectives.
The dataset provides a diverse set of sensor modalities to support neural reconstruction tasks. High-resolution RGB imagery is captured using a simulated camera at 4032$\times$3024 resolution with a 75 horizontal field of view. To encode geometric priors, ClaraVid includes metric depth maps and scene-level sparse point clouds (color, semantic, instance) sampled at three resolutions (30 m, 100 m, and 200 m), simulating LiDAR scans. For semantic understanding, panoptic segmentation masks provide instance-aware annotations for static structures such as buildings and dynamic entities including vehicles and pedestrians, with a semantic vocabulary of 18 classes. A key feature is the inclusion of dynamic object masks, which isolate moving entities, enabling motion-aware processing—an essential component for tasks requiring temporal coherence, such as NeRFs and dynamic scene reconstruction, to the best of our knowledge, feature unavailable in prior works. These multimodal annotations position ClaraVid as a comprehensive dataset for both static and dynamic aerial perception challenges. We present the delentropic scene profiles of all scenes in Figure 3, for a better understanding of the proposed dataset and provide additional samples in Section 10.
4.1 Development Environment
ClaraVid is developed using Unreal Engine, and leveraging the Carla[53] ecosystem for sensor fidelity and dynamic agent behavior, ensuring diverse yet reproducible traffic patterns. The dataset spans 5 distinct environments. For urban, rural, and highway scenarios, we enhance Carla’s original towns with photorealistic assets, refining textures, lighting, and object placement to elevate scene fidelity beyond default synthetic environments. The natural landscape is procedurally synthesized using real-world elevation [54] and road network priors from OSM[55], ensuring topographic realism, while the high-density urban environment is constructed from professionally designed assets that introduce architectural complexity. Across all scenes, procedural generation enriches foliage and ground cover—introducing dynamic trees, grass, and natural elements—to counteract the overly uniform, planar compositions found in prior synthetic datasets ([12, 13]. Manual refinements ensure spatial coherence and realism, creating environments that balance controlled benchmarking with the visual and structural richness needed for real-world applicability.
5 Experiments
We conduct a structured evaluation of ClaraVid to assess reconstruction and perception tasks across varying scene complexities and viewpoints. First, we benchmark the reconstruction accuracy of multiple neural reconstruction models on complete scene sets, analyzing their performance through delentropic scene profiling to quantify the impact of scene complexity on representational capacity. Second, we perform an ablation study on the influence of training viewpoints on multi-modal tasks, including novel view synthesis, semantic segmentation, and depth estimation, to evaluate each model’s adaptability to unseen viewpoints. Finally, we validate the dataset’s utility for downstream tasks by analyzing semantic segmentation results, providing a quantitative assessment of the synthetic-to-real domain gap.
Table 2: Novel view synthesis evaluation on ClaraVid. We compare multiple neural reconstruction methods under varying DSP, characterized by the mean ($\mu$) and shape parameters ($\alpha,\beta$). Higher $\mu$ values strongly correlate with reduced PSNR/SSIM ($\rho\approx-0.68$) and increased LPIPS ($\rho\approx+0.67$), indicating that greater scene complexity degrades reconstruction quality. $\alpha$ shows weak direct correlations, while a higher $\beta$, which skews the distribution toward simpler views, partially offsets performance loss by exhibiting a positive correlation with PSNR/SSIM and a negative correlation with LPIPS ($\rho\approx-0.3$).
Scene DSP Instant-NGP[36] TensoRF [56] Nerfacto[57] ZipNeRF[58] Gaussian Splatting[7]
$\mu$ $\alpha$ $\beta$ PSNR ↑ SSIM ↑ LPIPS ↓ PSNR ↑ SSIM ↑ LPIPS ↓ PSNR ↑ SSIM ↑ LPIPS ↓ PSNR ↑ SSIM ↑ LPIPS ↓ PSNR ↑ SSIM ↑ LPIPS ↓
Rural #1 4.18 4.50 1.65 18.27 0.3171 0.8492 18.11 0.2898 0.9204 18.36 0.2981 0.9022 18.56 0.4462 0.8818 19.97 0.5351 0.6019
Rural #2 3.98 3.08 2.39 20.75 0.3769 0.7192 20.61 0.3525 0.7890 21.29 0.3721 0.8145 24.54 0.5010 0.6645 22.59 0.5402 0.5727
Urban #1 3.67 10.36 4.47 22.25 0.5507 0.4829 21.96 0.5230 0.5271 23.24 0.5680 0.5321 26.39 0.7451 0.3614 26.37 0.7507 0.2981
Highway #1 3.98 10.76 2.87 20.32 0.3165 0.9148 20.41 0.3240 0.9043 20.96 0.3548 0.8862 21.73 0.4792 0.7960 22.03 0.4769 0.6814
Highway #2 3.55 1.66 1.03 22.97 0.5174 0.6389 22.78 0.4944 0.6887 22.66 0.5238 0.6668 24.51 0.6964 0.4609 24.61 0.6504 0.4858
Urban #2 3.29 3.29 1.74 23.85 0.5734 0.4925 23.66 0.5511 0.5438 24.24 0.5798 0.5485 25.42 0.6945 0.3443 26.68 0.7125 0.3658
Nature 4.12 3.94 3.12 21.40 0.3993 0.7057 21.02 0.3338 0.8213 21.60 0.3772 0.7705 24.55 0.5184 0.5829 22.94 0.5586 0.5515
Urban High 3.84 5.41 4.35 21.26 0.4878 0.5150 20.75 0.4465 0.5564 22.39 0.4797 0.5893 24.30 0.6920 0.4049 24.19 0.6211 0.4324
Average 3.97 5.37 2.70 21.38 0.4424 0.6648 21.16 0.4144 0.7189 21.84 0.4442 0.7138 23.75 0.5966 0.5622 23.67 0.6057 0.4987
Figure 4: Correlation between Delentropy and Reconstruction Score. Delentropy strongly correlates (Pearson) with reconstruction difficulty: higher delentropy corresponds directly to increased perceptual errors (LPIPS) and inversely to texture(PSNR) and geometric metrics(SSIM). The results are computed at image level aggregated across test splits.
Figure 5: Viewpoint Generalization in Neural Reconstruction. Comparison of Nerfacto and 3DGS across RGB, semantics, and depth. Trained on 90°||||H, Nerfacto generalizes better to unseen views (45°||||L, 55°||||M), while 3DGS excels in depth but degrades under shifts. Semantic accuracy follows reconstruction trends, with 3DGS performing well at training views but losing consistency under extrapolation.
5.1 Reconstruction Performance
We evaluate the reconstruction performance of various neural reconstruction models—Nerfacto [57], TensoRF [56], Instant-NGP [36], ZipNeRF [58], and Gaussian Splatting (3DGS) [9]—on the ClaraVid dataset, covering diverse scenarios characterized by distinct $DSP_{s}$. Performance is evaluated using PSNR, SSIM, and LPIPS, with results summarized in Table 2. A clear trend emerges where increased scene complexity, indicated by higher DSP values, leads to a consistent decline in reconstruction fidelity. This is reflected in the strong negative correlation between PSNR/SSIM and DSP mean values ($\rho\approx-0.68$), alongside a positive correlation with LPIPS ($\rho\approx+0.67$), signifying heightened perceptual error in more intricate scenes. While Gaussian Splatting and ZipNeRF demonstrate superior accuracy in structured urban environments, their advantage diminishes in more complex rural and natural landscapes. Interestingly, some exceptions arise. Namely, in natural scenes with substantial structural variation, NeRF-based methods maintain relatively higher PSNR scores, suggesting an ability to capture dominant large-scale structures effectively. However, these same models exhibit weaker perceptual quality, as reflected in their SSIM and LPIPS scores, indicating limitations in preserving fine details. A higher $\beta$ appears to mitigate reconstruction degradation, suggesting that scenes with simpler images are less susceptible to performance drops. Quantitative correlations between delentropy and reconstruction error across datasets further validate delentropy as a robust predictor of reconstruction difficulty (see Figure 4). We recognize the potential limitation of evaluating only synthetic data and explore its real-world applicability, detailed in Section  8.
Table 3: Performance Across Varying Viewpoints . Evaluation on appearance, segmentation, and depth synthesis across various viewpoints: $45^{\circ}|\text{{L}ow}$, $55^{\circ}|\text{{M}id}$, and $90^{\circ}|\text{{H}igh}$. (pitch||||altitude)
Nerfacto[57] Gaussian Splatting[7]
[…]
Appearance Synthesis
PSNR $\uparrow$
$45^{\circ}$ L 21.36 20.46 19.08 22.98 21.56 20.50
$55^{\circ}$ M 20.25 21.43 20.48 20.75 22.75 21.06
$90^{\circ}$ H 18.23 19.42 22.45 17.88 19.26 23.00
SSIM $\uparrow$
$45^{\circ}$ L 0.42 0.40 0.36 0.58 0.50 0.42
$55^{\circ}$ M 0.39 0.42 0.40 0.45 0.58 0.41
$90^{\circ}$ H 0.34 0.40 0.47 0.36 0.65 0.59
LPIPS $\downarrow$
$45^{\circ}$ L 0.75 0.74 0.76 0.55 0.61 0.78
$55^{\circ}$ M 0.77 0.73 0.74 0.59 0.53 0.62
$90^{\circ}$ H 0.79 0.71 0.68 0.81 0.73 0.50
Semantic Segmentation
mIoU $\uparrow$
$45^{\circ}$ L 44.82 44.14 40.78 56.53 55.11 53.50
$55^{\circ}$ M 43.12 46.57 45.74 53.34 59.59 56.50
$90^{\circ}$ H 36.95 41.64 51.08 36.68 40.82 58.54
Depth
Abs Rel $\downarrow$
$45^{\circ}$ L 0.059 0.042 0.066 0.025 0.017 0.010
$55^{\circ}$ M 0.042 0.043 0.027 0.023 0.016 0.009
$90^{\circ}$ H 0.054 0.039 0.022 0.028 0.019 0.010
RMSE $\downarrow$
$45^{\circ}$ L 12.34 10.64 14.22 4.67 3.37 2.36
$55^{\circ}$ M 11.47 9.09 6.55 4.61 3.07 1.93
$90^{\circ}$ H 14.20 8.95 4.68 5.41 3.53 2.08
$\boldsymbol{\delta<1.05}$uparrow
$45^{\circ}$ L 0.8193 0.8467 0.8317 0.881 0.92 0.95
$55^{\circ}$ M 0.8315 0.8584 0.9024 0.895 0.93 0.96
$90^{\circ}$ H 0.7886 0.8354 0.8980 0.871 0.91 0.95
5.2 Performance Across Varying Viewpoints
We evaluate the generalization capability of 2 neural reconstruction paradigms—NeRF based NeRF (Nerfacto) and particle-based Gaussian Splatting (3DGS) — in the context of aerial viewpoint variations. This experiment is motivated by real-world UAV-based mapping scenarios, where viewpoint interchangeability is essential (e.g., oblique images intended for perspective mapping may later be repurposed to generate nadir orthophotos). Each method is trained on images from a single altitude angle ($45^{\circ}$, $55^{\circ}$, or $90^{\circ}$) and subsequently tested on unseen viewpoints to assess robustness under realistic operational conditions. The quantitative results, detailed in Table 3, reveal distinct performance profiles across novel view synthesis, semantic segmentation (mIoU), and depth estimation (RMSE, AbsRel, and $\delta^{1.05}$). Nerfacto exhibits greater robustness to viewpoint variations, maintaining relatively consistent reconstruction quality across novel views, albeit with longer training times. Conversely, 3DGS achieves higher fidelity on familiar views and converges faster but degrades significantly under moderate viewpoint shifts, limiting its suitability for applications requiring broad viewpoint coverage. Semantic segmentation accuracy (mIoU) closely follows reconstruction quality trends, with 3DGS performing better at training viewpoints but losing accuracy under extrapolation, suggesting a strong dependence on appearance consistency. Depth estimation results, however, favor 3DGS, which consistently outperforms Nerfacto, indicating superior geometric reconstruction. We used a stricter depth accuracy threshold of $\delta<1.05$, as even minor inaccuracies can propagate into large errors at great distances ($>100m$). For visual comparison see Figure 5.
Table 4: Performance on semantic segmentation. Entropy refers to the normalized Shannon for measuring prediction confidence.
[…]
Oracle/Baseline
UAVid[24] UAVid 0.144 71.36
UDD[22] UDD 0.047 79.84
SkyScapes[30] SkyScapes 0.176 65.49
ClaraVid ClaraVid 0.095 82.51
Syndrone[12] Syndrone 0.051 86.36
Synthetic to Real
ClaraVid UAVid 0.213 57.50
Syndrone UAVid 0.235 52.49
ClaraVid UDD 0.196 58.22
Syndrone UDD 0.213 54.43
ClaraVid SkyScapes 0.290 47.52
Syndrone SkyScapes 0.310 42.75
5.3 Semantic Segmentation as downstream task
We assess ClaraVid’s effectiveness in downstream tasks by quantifying the synt-to-real domain gap in semantic segmentation. A segmentation model (see Section 9 for details) is trained on real [24, 22, 30] and synthetic datasets, with results shown in Table 4. Despite containing 4.5$\times$ fewer samples than Syndrone [12], ClaraVid achieves a smaller domain gap, demonstrating superior scene complexity and image fidelity. Models trained on ClaraVid consistently outperform those trained on Syndrone across all real datasets, achieving higher mIoU and lower prediction entropy. However, a residual semantic gap remains, with ClaraVid exhibiting an average gap of 17.82%, compared to 22.34% for Syndrone.
6 Conclusions
In this work, we introduced ClaraVid, a synthetic aerial dataset designed to bridge the gap in holistic scene reconstruction tailored for UAV-based scene understanding. We further proposed the Delentropic Scene Profile, a novel complexity metric rooted in differential entropy, which quantifies scene difficulty and exhibits strong correlations with reconstruction performance. Our benchmarking framework systematically evaluates neural reconstruction techniques, revealing how DSP serves as a reliable predictor of reconstruction accuracy, informing method selection and optimization strategies. Through extensive experimentation, we demonstrate that ClaraVid not only surpasses existing synthetic datasets in realism and annotation quality but also generalizes better to real-world UAV imagery, challenging the conventional notion that dataset size alone dictates generalization. By releasing ClaraVid, we aim to advance UAV-based perception and improve benchmarking for semantic and geometric reconstruction.
Supplementary Material
7 Delentropic Scene Profile Real Data
To evaluate the effectiveness of the Delentropic Scene Profile (DSP) in real-world environments beyond synthetic datasets, we follow the experimental framework of CityGaussian [45] and assess DSP on urban mapping datasets, including UrbanScene3D [11] and Mill19 [46]. These datasets feature diverse urban landscapes with varying degrees of structural complexity, allowing us to analyze the correlation between delentropy and reconstruction performance under real-world conditions. By computing the delentropy profile for each scene and aggregating the results, we establish a direct comparison with the experimental numbers reported in these studies. As shown in  Figure 6, the scenes exhibit moderate complexity, with the residential scene in particular displaying a DSP profile comparable to the Urban High category in the Claravid dataset. Additionally, we examine large-scale NeRF-based methods, such as Mega-NeRF [46], Switch-NeRF [59], CityGaussian [45], Gaussian Splatting [7], and GN-NeRF [60]. Our findings, present in Figure 7, reveal that higher DSP values are consistently associated with increased reconstruction errors, reinforcing delentropy’s predictive capability in real-world scenarios. These results establish DSP as a robust metric for quantifying scene complexity and estimating reconstruction difficulty.
Figure 6: Delentropy profiles for real-world mapping datasets. The DSP indicates medium complexity, with variations reflecting structural differences across scenes. The profile is computed across all scene images.
Figure 7: Reconstruction quality vs. scene complexity. The relationship between scene complexity, measured by the $\mu$ mean delentropy, and reconstruction fidelity (PSNR) across different neural reconstruction methods. A negative correlation is observed, where higher delentropy values correspond to lower PSNR scores, indicating increased reconstruction difficulty.
8 Experimental setup
8.1 Reconstruction Performance
Throughout our experiments, we utilize the Nerfstudio framework[57] to systematically benchmark various neural reconstruction methods. For Nerfacto[57], we adopt the big configuration. TensoRF[56] is configured with an increased grid resolution of $500^{3}$. InstantNGP[36] employs 16 grid levels, a maximum resolution of 8192, and a hash map size of $2^{21}$. Zip-NeRF[58] is used in its default configuration without modifications. Gaussian Splatting[7] is initialized from the dataset’s scene-level point cloud with a spatial resolution of 100 cm, restricting point splitting beyond 3 million points to balance memory efficiency and reconstruction accuracy. Training durations vary across models, ranging from 60 to 720 minutes on an NVIDIA A6000 GPU, depending on architectural complexity. Across all experiments, we leverage FP16 training to optimize memory usage. The reported reconstruction and segmentation results correspond to a half-resolution output relative to the original image dimensions, maintaining a trade-off between computational feasibility and fidelity in large-scale aerial scene reconstruction.
8.2 Performance Across Varying Viewpoints
Both Nerfacto and Gaussian Splatting use the configuration from Section 8.2. For semantic learning, Nerfacto incorporates a jointly trained segmentation field, while Gaussian Splatting learns a separate set of 32-width semantic parameters—trained jointly but independent of geometry or positional encodings. Additionally, a 2-layer CNN classifier refines segmentation on the Gaussian splats. We compute per scene class weighting to balance semantic segmentation classes.
Figure 8: Semantic Segmentation Pixel Label Distribution The class distribution exhibits a long tail characteristic for aerial scenarios.
9 Semantic Segmentation
We train a DPT[61] model with a DinoV2-L[62] backbone on both synthetic and real-world UAV datasets. The backbone is kept frozen while only the decoder is optimized. Our training protocol is standardized across datasets, with the number of steps adjusted proportionally to the dataset size—ranging from 5,000 to 30,000—to prevent overfitting. We use each dataset’s native resolution and perform random crops of 630×630 pixels. Additionally, we uniformly apply data augmentations such as color adjustments, lighting variations, rotations, and distortions. Finally, the semantic palette is adapted to the real dataset following the [12] protocol.
10 Claravid
In this section, we present additional insights into Claravid.
10.1 Semantic Complexity
The semantic complexity of Claravid is a direct consequence of its enriched class taxonomy, designed to capture the nuanced structures present in aerial urban and rural environments. By extending the conventional label set to include wire, a class that encapsulates thin linear structures, and green energy, which aggregates solar panels and renewable energy infrastructure, the dataset reflects the intricate composition of real-world landscapes. Furthermore, the pole category has been redefined to encompass slender metallic structures such as communication towers and high-voltage power lines, thus improving its generalization for fine-grained segmentation tasks. Additionally, the introduction of urban props and construction props provides a finer semantic partitioning of the environment, accounting for human-centric elements in residential areas and industrial material clusters, respectively. The inclusion of a dedicated tent class further enhances the granularity of temporary and semi-permanent structures. This expanded semantic schema not only increases the dataset’s diversity but also contributes to its delentropy-based complexity profiling, allowing for a more rigorous quantification of scene heterogeneity in both structural and semantic dimensions. We present the label class distribution in Figure 8 and complementary, the depth distribution across the entire dataset in Figure 9.
Figure 9: Dataset Depth Distribution The dataset’s depth values exhibit a broad distribution with prominent spikes from nadir imagery, where ground-level elements dominate. While depths range from 0–1000 m, the 99th percentile is 192[m].
10.2 Rendering Pipeline Configuration
To achieve high-fidelity rendering of large-scale aerial scenes in Unreal Engine 4 (UE4), we implement targeted modifications to the rendering pipeline, prioritizing fine detail preservation and visibility consistency at extended distances. To mitigate the disappearance of distant objects and enhance detail clarity, we increase the rendering resolution by setting r.ScreenPercentage=300, effectively supersampling the scene at 3 times the native resolution. This adjustment minimizes aliasing artifacts—particularly pronounced in oblique aerial perspectives—and ensures that small-scale features remain discernible, albeit at a higher computational cost justified by the resulting visual fidelity. Shadow integrity for distant fine elements, such as foliage and thin structures, is preserved by reducing r.Shadow.MinRadius to 0.001, which enables shadow casting for geometrically narrow features that would otherwise be lost. Shadow map resolution is also increased to 2048×2048 for objects considered having fine details, maintaining sharp shadow edges across expansive terrains. Geometric fidelity is ensured by disabling level-of-detail (LOD) transitions, preventing mesh simplification that degrades structural complexity at greater distances or altitudes, a choice prioritizing rendering quality over real-time performance scalability. Similarly, we disable distance-based object culling to maintain persistent visibility of all scene elements, eliminating sudden visibility discontinuities that disrupt spatial and temporal coherence critical for synthetic aerial datasets. We present a visual overview of the dataset in Section 10.
Figure 10: Claravid Overview. Representative images across 5 different scenes that were either created or improved. From top to bottom: Urban Dense, Nature, Rural, Urban, Highway.